[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Oct 17, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Oct 17 21:02:32 UTC 2016


HSU Shift Log for Mon Oct 17 2016 17:02:26 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: ATCF was sluggish with Nicole due to large files resulting  
from 55 advisories; EPAC workstation locked up for the 15Z advisory;  
rebooted all ATCF systems, which helped speed up graphics.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161017/84cebcec/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list